OpenStudy (anonymous):

(AP Physics) Can you please tell me how to figure this out? How do we find the initial velocity of a projectile that travels a distance of 52.5 meters and reaches a maximum height of 30 meters. How can I figure that out?

OpenStudy (anonymous):

x= Vicos(theta)*t y max = Visin(theta)*t/2 -1/2*g*(t/2)^2

OpenStudy (anonymous):

Sorry, you need a 3rd eqn. : 0 = Visin(theta)*t -1/2*g*t^2

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

I don't have " t "

OpenStudy (anonymous):

nope. but you have 3 equations with 3 unknowns. you know what to do:)

OpenStudy (anonymous):

ok. And one more question.

OpenStudy (anonymous):

looking at these eqn.s, I see a little strategy for solving quickly...

OpenStudy (anonymous):

ok. how? And thatnks for the help,

OpenStudy (anonymous):

x= Vxi*t (no acceleration) Vyf^2 = Viy^2 -2gh (solve at max height: Vyf =0) use Vxi and Vyi : Vi =sqrt(Vxi^2+Vyi^2)

OpenStudy (anonymous):

x =52.5 and h=30

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

I was solving and realized that they simplify into more compact equations...

OpenStudy (anonymous):

Yes. It makes a lot of sense now. Thank you. I appreciate the help.

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

oh yeah. And one more question. Since the projectile traveled a full length of 52.5, should I use 52.5 or just half of that?